A square building of side x meters is extended by 15m on one side and 10m on the other side to form a rectangle.
Express the new area of the rectangle as the product of two binomials

Answer by stanbon(75887)

A square building of side x meters is extended by 15m on one side and 10m on the other side to form a rectangle.
Express the new area of the rectangle as the product of two binomials
-------
New area = (x+15)(x+10) = x^2+25x+150
--------------------------------------
Cheers,
Stan H.
==========